$SPAD/src/input set.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool set.output
)set message test on
)set message auto off
)clear all
--S 1 of 20
s := set [x**2-1, y**2-1, z**2-1]
--R
--R
--R          2      2      2
--R   (1)  {x  - 1,y  - 1,z  - 1}
--R                                               Type: Set(Polynomial(Integer))
--E 1
--S 2 of 20
t := set [x**i - i+1 for i in 2..10 | prime? i]
--R
--R
--R          2      3      5      7
--R   (2)  {x  - 1,x  - 2,x  - 4,x  - 6}
--R                                               Type: Set(Polynomial(Integer))
--E 2
--S 3 of 20
i := intersect(s,t)
--R
--R
--R          2
--R   (3)  {x  - 1}
--R                                               Type: Set(Polynomial(Integer))
--E 3
--S 4 of 20
u := union(s,t)
--R
--R
--R          2      3      5      7      2      2
--R   (4)  {x  - 1,x  - 2,x  - 4,x  - 6,y  - 1,z  - 1}
--R                                               Type: Set(Polynomial(Integer))
--E 4

--S 5 of 20
difference(s,t)
--R
--R
--R          2      2
--R   (5)  {y  - 1,z  - 1}
--R                                               Type: Set(Polynomial(Integer))
--E 5
--S 6 of 20
symmetricDifference(s,t)
--R
--R
--R          3      5      7      2      2
--R   (6)  {x  - 2,x  - 4,x  - 6,y  - 1,z  - 1}
--R                                               Type: Set(Polynomial(Integer))
--E 6
--S 7 of 20
member?(y, s)
--R
--R
--R   (7)  false
--R                                                                Type: Boolean
--E 7
--S 8 of 20
member?((y+1)*(y-1), s)
--R
--R
--R   (8)  true
--R                                                                Type: Boolean
--E 8
--S 9 of 20
subset?(i, s)
--R
--R
--R   (9)  true
--R                                                                Type: Boolean
--E 9
--S 10 of 20
subset?(u, s)
--R
--R
--R   (10)  false
--R                                                                Type: Boolean
--E 10

--S 11 of 20
gs := set [g for i in 1..11 | primitive?(g := i::PF 11)]
--R
--R
--R   (11)  {2,6,7,8}
--R                                                    Type: Set(PrimeField(11))
--E 11
--S 12 of 20
complement gs
--R
--R
--R   (12)  {1,3,4,5,9,10,0}
--R                                                    Type: Set(PrimeField(11))
--E 12
--S 13 of 20
a := set [i**2 for i in 1..5]
--R
--R
--R   (13)  {1,4,9,16,25}
--R                                                   Type: Set(PositiveInteger)
--E 13
--S 14 of 20
insert!(32, a)
--R
--R
--R   (14)  {1,4,9,16,25,32}
--R                                                   Type: Set(PositiveInteger)
--E 14
--S 15 of 20
remove!(25, a)
--R
--R
--R   (15)  {1,4,9,16,32}
--R                                                   Type: Set(PositiveInteger)
--E 15
--S 16 of 20
a
--R
--R
--R   (16)  {1,4,9,16,32}
--R                                                   Type: Set(PositiveInteger)
--E 16
--S 17 of 20

b := b0 := set [i**2 for i in 1..5]
--R
--R
--R   (17)  {1,4,9,16,25}
--R                                                   Type: Set(PositiveInteger)
--E 17
--S 18 of 20
b := union(b, {32})
--R
--R
--R   (18)  {1,4,9,16,25,32}
--R                                                   Type: Set(PositiveInteger)
--E 18
--S 19 of 20
b := difference(b, {25})
--R
--R
--R   (19)  {1,4,9,16,32}
--R                                                   Type: Set(PositiveInteger)
--E 19
--S 20 of 20
b0
--R
--R
--R   (20)  {1,4,9,16,25}
--R                                                   Type: Set(PositiveInteger)
--E 20
)spool
)lisp (bye)

———-

References^{[1] nothing}

7